Solve the system of equations by substitution: y=x to the second power+2x-2
y=3x+4

Answer by checkley71(8403)

Y=X^2+2X-2 NOW SUBSTITUTE (X^2+2X-2) FOR Y IN THE SECOND EQUATION & SOLVE FOR X.
X^2+2X-2=3X+4
X^2+2X-3X-2-4=0
X^2-X-6=0
(X-3)(X+2)=0
X-3=0
X=3 ANSWER.
X+2=0
X=-2 ANSWER.
Y=3*3+4
Y=9+4
Y=13 ANSWER.
Y=3*-2+4
Y=-6+4
Y=-2 ANSWER.